Homework Assignment #9 (Due date: May 18th, 2016) Chapter 9 -The Nuclous, Radioactivity, and Nuclear Medicine (1) (2) (3) (4) What are the major differences How do nuclear reactions and chemical reaciions differ? Rank the three major types between alpha and beta p articles? (alpha, beta, and gamma) of radiation in order of size, speed, and penetrating power. (4) write a nuclear equation to represent cobalt-60 decaying to nickel-60 plus a beta particle and a Write a nuclear equation t gamma ray (5) Complete the following nuclear equations 140 (6) Would you predict oxygen-20 to be stable? Explain your reasoning. A patient receives 9.0 ng of a radioisotope with a half-like of 12 h. How much will remain in the body after 2.0 days, assuming that radioactive decay is the only path for removal of the isotope from the body? (7) bbey2 0 days, assuming that radioavdecayis the only pathfor removal of the isotope from the (8) Define each of the following units: (a) curie (b) roentgen X-ray technicians often wear badges containing photographic film. How is this film used to indicate exposure to X-ray? (9)
Answer #1

Post other question in different post

Q1) Major difference between alpha and beta particle

Alpha beta decay will cause tha decrease of 2 in the atomic number and 4 in the mass number whereas beta decay will doesn't affect the atomic number and will only affect the mass number

Alpha beta decay is a slow process whereas beta particle decay is a very fast process because of movement of fast moving electrons

Alpha particles deflect under the effect of magnetic field because of their high weight but the beta particle didn't get deflected

Q2)

Nuclear reaction involve the change in nucleus atom because of difference in either of atomic number and mass number whereas there is only rearrangement in the case of chemical reactions

Q6) The oxygen 20 won't be stable

Number of protons = 8

Number of neutrons = 20 - 8 = 12

Since the ration of neutrons to protons is equal to 1.5, which is very high and it must be close to 1 in order for the atom to be stabel

Q7) ln(Ao/At) = kt

ln(Ao/At) = ln2/12 & 48

Ao/At =16

At = Ao/16 = 9/16 mg = 0.5625 mg